OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17 A. x + 3(–2y + 2) = 17 B. (2y + 2) + 3y = 17 C. x + 3(2y + 2) = 17 D. (–2y + 2) + 3y = 17

OpenStudy (anonymous):

B

OpenStudy (anonymous):

just put x = 2y + 2 from 1st eqn in the place of x in 2nd eqn

OpenStudy (anonymous):

correct its B x – 2y = 2 x=2+2y sub this to the other eq x + 3y = 17 (2+2y)+3y=17

OpenStudy (anonymous):

B. (2y + 2) + 3y = 17 x – 2y = 2 x + 3y = 17 ----------- solve eqn x – 2y = 2 for x. we get, x=2y+2 plug this value of x in eqn x + 3y = 17 to get (2y+2) + 3y=17